# Seeing Structural Failure Before it Happens: An Image-Based Physics-Informed Neural Network (PINN) for Spaghetti Bridge Load Prediction

### v2

@@ I Introduction
Physics Informed Neural Networks (PINNs) have emerged as a powerful class of models that integrate known physical laws into the training process of neural networks [1]. By embedding physics directly into the loss function, PINNs ensure that the learned solutions not only fit the data but also satisfy governing physical relationships. This hybrid approach has demonstrated superior generalization and robustness compared to purely data-driven methods, particularly in settings with limited or noisy data. PINNs have been successfully applied in a range of scientific and engineering domains, including fluid dynamics [2], biomedical engineering [3], geophysics [4], and heat transfer problems [5], where they have consistently outperformed conventional neural networks in producing physically consistent and interpretable results.
In structural engineering, traditional analysis techniques such as those based on finite element methods, have long provided reliable results, but often at the cost of high computational expense and a reliance on detailed, sometimes hard-to-measure input parameters [6, 7]. As structural systems grow in complexity and scale, there is a growing need for predictive models that are not only computationally efficient but also capable of generalizing across varying configurations [8, 9]. While machine learning (ML) methods offer promising alternatives, their purely data-driven nature can lead to predictions that overlook essential physical constraints and may be physically implausible [10]. In this context, PINNs offer a compelling middle ground that blends data-driven flexibility with embedded domain knowledge to produce results that are both accurate and physically meaningful [11, 12].
To explore the practical application of PINNs within the context of structural engineering, we focus on spaghetti bridges which is a simplified yet instructive case. These small scale models serve as effective proxies for investigating the behavior of larger structural systems due to which they are widely used in educational settings to demonstrate fundamental principles of structural mechanics [13]. Despite their simplicity, spaghetti bridges exhibit key mechanical behaviors found in full scale structures, including load transfer, stress distribution, and characteristic failure modes [14].
@@ II Related Work
A physics-informed neural network (PINN) framework is proposed to predict the dynamic responses of bridges under moving loads [15]. The partial differential equations and boundary conditions are integrated into the training process of the model that achieves accurate spatiotemporal stress and displacement predictions. Consequently, the approach reduces reliance on dense sensor networks and shows high accuracy across various load scenarios.
@@
A physics-informed neural network (PINN)-based framework is propsed to identify bridge influence lines and dynamic loads from multiple vehicles [16]. Using sparse sensors data, their model integrates bridge mechanics into the training process of PINN to predict structural responses. The deployment of PINNs achieve improved results by accurately estimating both the location and magnitude of loads compared to the conventional ML-based approaches.
@@
The PINN-based structural health monitoring is performed by modeling the behavior of Kirchhoff–Love plates that are used to represent bridge decks and thin-walled structures [17]. With limited and noisy data, the plate theory-induced learning process of PINN results in the successful detection of structural damage. Consequential, the achieved results demonstrate that PINNs can effectively capture physical behavior and damage locations without relying on conventional dense instrumentation.
@@
a physics-informed recurrent neural network (PI-RNN) is proposed for the reconstruction of dynamic displacement in bridges under moving loads [18]. The RNN architecture is enriched with the physical properties of the bridge by integrating their respective equations. As a result, the PI-RNN model accurately recovers full-field displacements from sparse acceleration data. The proposed method is shown to outperform conventional neural networks demonstrating strong potential for real-time structural health monitoring.
@@
Yet again, under moving loads on girder bridges, a PINN-based virtual sensing method is proposed to predict structural responses [19]. By embedding physical laws into the training process, the proposed model successfully reconstructs displacements and internal forces using limited sensor data. The approach significantly reduces dependence on dense instrumentation and provides accurate, real-time estimations for bridge monitoring.
@@
For predicting the response of highway bridges under dynamic loads due to moving vehicles, a PINN-based approach is proposed [20]. By incorporating governing equations into the learning framework, the model accurately simulates displacement and acceleration with minimal sensor input. The results demonstrate that PINNs can offer a reliable, physics-consistent alternative to conventional numerical simulations in bridge engineering.
@@
A novel PINN-based framework is proposed for identifying damage in multi-degree-of-freedom reinforced concrete bridge piers subjected to nonlinear dynamic loading [21]. The model integrates nonlinear structural dynamics into the training process of PINN, enabling detection of damage parameters without requiring full-scale physical sensors.
@@
Most of these studies focus on large structures and require detailed information or many sensors. One related study by Patel and Rodriguez [22] used traditional neural networks to predict the failure load of small-scale educational bridges. Although their results are encouraging, their approach relied on manual geometric measurements and did not include any physics-based constraints in the learning process. In our work, we take a different path by introducing physics informed learning to spaghetti bridge modeling. To the best of our knowledge, this is the first study that applies a physics informed neural network to these types of bridges. We believe this approach offers better reliability because it not only learns from the data but also respects the underlying physical rules of structural behavior. To achieve this, we propose a new model called the Physics Informed Kolmogorov Arnold Network, which combines mathematical learning with structural physics. Inspired by StrucNet [23], we also present a computer vision method that detects and extracts bridge parameters directly from images, making the system more practical and easier to use. This combination allows us to show that accurate predictions are possible even when the dataset is small, which is often the case in educational setups.
@@ III-A Initial Data Collection
[…]
@@ VI-A Base Neural Network Architecture
The core architecture of our model consists of:

### v4

@@ I Introduction
Physics Informed Neural Networks (PINNs) have emerged as a powerful class of models that integrate known physical laws into the training process of neural networks [raissi2019physics]. By embedding physics directly into the loss function, PINNs ensure that the learned solutions not only fit the data but also satisfy governing physical relationships. This hybrid approach has demonstrated superior generalization and robustness compared to purely data-driven methods, particularly in settings with limited or noisy data. PINNs have been successfully applied in a range of scientific and engineering domains, including fluid dynamics [ren2024physics], biomedical engineering [kissas2020machine], geophysics [secci2024physics], and heat transfer problems [jalili2024physics], where they have consistently outperformed conventional neural networks in producing physically consistent and interpretable results.
In structural engineering, traditional analysis techniques such as those based on finite element methods, have long provided reliable results, but often at the cost of high computational expense and a reliance on detailed, sometimes hard-to-measure input parameters [liu2025investigation, dilsiz2025field]. As structural systems grow in complexity and scale, there is a growing need for predictive models that are not only computationally efficient but also capable of generalizing across varying configurations [zhang2025full, das2025hybrid]. While machine learning (ML) methods offer promising alternatives, their purely data-driven nature can lead to predictions that overlook essential physical constraints and may be physically implausible [urban2025unveiling]. In this context, PINNs offer a compelling middle ground that blends data-driven flexibility with embedded domain knowledge to produce results that are both accurate and physically meaningful [chew2025physics, wang2025physics].
To explore the practical application of PINNs within the context of structural engineering, we focus on spaghetti bridges which is a simplified yet instructive case. These small scale models serve as effective proxies for investigating the behavior of larger structural systems due to which they are widely used in educational settings to demonstrate fundamental principles of structural mechanics [groth2018framework]. Despite their simplicity, spaghetti bridges exhibit key mechanical behaviors found in full scale structures, including load transfer, stress distribution, and characteristic failure modes [mahendran1995project].
@@ II Related Work
A physics-informed neural network (PINN) framework is proposed to predict the dynamic responses of bridges under moving loads [li2025moving]. The partial differential equations and boundary conditions are integrated into the training process of the model that achieves accurate spatiotemporal stress and displacement predictions. Consequently, the approach reduces reliance on dense sensor networks and shows high accuracy across various load scenarios.
@@
A physics-informed neural network (PINN)-based framework is propsed to identify bridge influence lines and dynamic loads from multiple vehicles [li2025identification]. Using sparse sensors data, their model integrates bridge mechanics into the training process of PINN to predict structural responses. The deployment of PINNs achieve improved results by accurately estimating both the location and magnitude of loads compared to the conventional ML-based approaches.
@@
The PINN-based structural health monitoring is performed by modeling the behavior of Kirchhoff–Love plates that are used to represent bridge decks and thin-walled structures [al2024physics]. With limited and noisy data, the plate theory-induced learning process of PINN results in the successful detection of structural damage. Consequential, the achieved results demonstrate that PINNs can effectively capture physical behavior and damage locations without relying on conventional dense instrumentation.
@@
a physics-informed recurrent neural network (PI-RNN) is proposed for the reconstruction of dynamic displacement in bridges under moving loads [tao2025dynamic]. The RNN architecture is enriched with the physical properties of the bridge by integrating their respective equations. As a result, the PI-RNN model accurately recovers full-field displacements from sparse acceleration data. The proposed method is shown to outperform conventional neural networks demonstrating strong potential for real-time structural health monitoring.
@@
Yet again, under moving loads on girder bridges, a PINN-based virtual sensing method is proposed to predict structural responses [al2025developing]. By embedding physical laws into the training process, the proposed model successfully reconstructs displacements and internal forces using limited sensor data. The approach significantly reduces dependence on dense instrumentation and provides accurate, real-time estimations for bridge monitoring.
@@
For predicting the response of highway bridges under dynamic loads due to moving vehicles, a PINN-based approach is proposed [yousefpour2025physics]. By incorporating governing equations into the learning framework, the model accurately simulates displacement and acceleration with minimal sensor input. The results demonstrate that PINNs can offer a reliable, physics-consistent alternative to conventional numerical simulations in bridge engineering.
@@
A novel PINN-based framework is proposed for identifying damage in multi-degree-of-freedom reinforced concrete bridge piers subjected to nonlinear dynamic loading [yamaguchi2024physics]. The model integrates nonlinear structural dynamics into the training process of PINN, enabling detection of damage parameters without requiring full-scale physical sensors.
@@
Most of these studies focus on large structures and require detailed information or many sensors. One related study by Patel and Rodriguez [b4] used traditional neural networks to predict the failure load of small-scale educational bridges. Although their results are encouraging, their approach relied on manual geometric measurements and did not include any physics-based constraints in the learning process. In our work, we take a different path by introducing physics informed learning to spaghetti bridge modeling. To the best of our knowledge, this is the first study that applies a physics informed neural network to these types of bridges. We believe this approach offers better reliability because it not only learns from the data but also respects the underlying physical rules of structural behavior. To achieve this, we propose a new model called the Physics Informed Kolmogorov Arnold Network, which combines mathematical learning with structural physics. Inspired by StrucNet [b2], we also present a computer vision method that detects and extracts bridge parameters directly from images, making the system more practical and easier to use. This combination allows us to show that accurate predictions are possible even when the dataset is small, which is often the case in educational setups.
@@ III Dataset Preparation
Figure 1 presents the complete workflow of our proposed approach, from initial data collection through model evaluation and result visualization.
Figure 1: Overview of the complete methodology pipeline: (1) Initial dataset of 15 real bridges with measured geometric and material properties, (2) Data augmentation through parameter variation, noise addition, and physics-consistent transformations to generate 100 samples, (3) Preprocessing including standardization and 80-20 train-test split, (4) Training of PINN and PIKAN models with physics-informed loss functions, (5) Predictions on test set, (6) Evaluation metrics achieving MAE of 10.50g, RMSE of 13.37g, and R² of 0.9603, (7) Result visualization and comparison including physics loss component analysis.
@@ III-A Initial Data Collection
Each bridge was constructed using standard dry spaghetti strands and carefully measured to ensure accurate parameter recording. The weight was measured using a precision digital scale with $\pm$0.1g accuracy.
@@ VI-A Base Neural Network Architecture
Figure 3: Base neural network architecture used for weight prediction. The model consists of three fully connected hidden layers with ReLU activation, Batch Normalization, and Dropout regularization.
The core architecture of our model, illustrated in Fig. 3, consists of:
@@ VI-B Physics-Informed Loss Function
As illustrated in Figure 4, the physics-informed loss function integrates both data-driven and physics-driven components to guide the training of the PIKAN model.
Figure 4: Schematic of the Physics-Informed Loss Function used in the PIKAN model, showing the combination of data-driven and physics-driven loss components.
@@ VII-B1 TruncatedPolynomialLayer
As shown in Figure 5, the Truncated Polynomial Layer expands input features into higher-order polynomial terms and pairwise interactions.
Figure 5: Illustration of the Truncated Polynomial Layer used in the PIKAN model. The layer expands input features into higher-order polynomial terms and pairwise interactions up to a fixed degree.